# Towards Optimizing the Expected Performance of Sampling-Based Quantum-Inspired Algorithms

### v2

@@ 1 Introduction
Quantum algorithms have intrigued researchers for years with the potential of achieving exponential speedup compared to classical algorithms [10, 39, 36, 31, 24, 8]. These algorithms leverage the principles of quantum mechanics to perform computations in parallel, potentially solving certain problems much more efficiently than classical algorithms ever could.
However, contrary to prevailing belief, it was recently shown that under certain conditions—when data is structured in a particular way—some quantum machine learning (QML) algorithms do not exhibit exponential speedup over classical ones [15, 40, 41, 7]. In other words, the exponential speedup of quantum algorithms relied on the assumption that preparing the quantum state encoding of a given classical input could be achieved in sublinear time relative to the input size. This fact was first demonstrated by explicitly constructing a classical recommendation algorithm inspired by its quantum counterpart [40]. If classical data is preprocessed in a proper manner, a dequantized QML algorithm can be designed to run only polynomially slower than the quantum version. Specifically, the data must be prepared in a tree structure to enable efficient sampling comparable to quantum algorithms.
Since the introduction of the quantum-inspired recommendation system, several quantum algorithms have been successfully dequantized [15, 41, 7], highlighting both the limitation and potential of quantum computing. Under certain conditions, these algorithms are expected to achieve much better performance than known theoretical bounds, especially in extremely high-dimensional problems [1]. Therefore, it is instructive to characterize the best possible performance of these quantum-inspired algorithms for precise comparison and further development.
@@
Direct Fidelity Estimation (DFE) [13], a technique developed in advance of quantum-inspired algorithms, is another practical application of sampling-based inner product estimation. In Section 5, we show that for some target states, we can tighten the upper bound on the number of required measurements to perform DFE by sampling a Pauli operator based on the $L^{1}$ norm distribution for the characteristic function.
@@ 2.1 Nomenclature
Consider $x\in\mathbb{R}^{n}$ and $A\in\mathbb{R}^{m\times n}$. $||x||$ and $||x||_{p}$ are the $L^{2}$ norm and $L^{p}$ norm of $x$, respectively. $||A||_{F}$ is the Frobenius norm of $A$. $A^{(j)}$ is the $j$-th column of $A$. $\tilde{A}$ is a vector in $\mathbb{R}^{n}$ such that $\tilde{A}_{j}=||A^{(j)}||$ and $\tilde{A}^{(p)}$ is a vector in $\mathbb{R}^{n}$ such that $\tilde{A}^{(p)}_{j}=||A^{(j)}||_{p}$. $\text{nnz}(x)$ ($\text{nnz}(A)$) is the number of nonzero elements in $x$ ($A$). $\mathcal{U}_{(a,b)}$ is the continuous uniform distribution over $(a,b)$. $[n]$ denotes the set $\{1,2,\ldots,n\}$.
@@ 2.2 $\text{SQ}_{2}$ data structure
$\mathcal{D}_{x}(i)=\frac{x_{i}^{2}}{||x||^{2}}.$
Hereafter, the notation $\mathcal{D}_{x}$ implies that $x$ is a nonzero vector. Definition 1 has a quantum analog, since $\mathcal{D}_{x}$ is the distribution corresponding to the amplitude encoding of $x$ [38]:
@@
Figure 2: BST data structures for (a) $x\in\mathbb{R}^{4}$ and (b) $A\in\mathbb{R}^{4\times 4}$ based on $L^{p}$ norm.
@@ 2.3 Inner product estimation with $\text{SQ}_{2}$
Given $\textnormal{SQ}_{2}(x)$ and $\textnormal{Q}(y)$ for $x,y\in\mathbb{R}^{n}$, the inner product $\langle x,y\rangle$ can be estimated to additive error $\epsilon||x||\cdot||y||$ with at least $1-\delta$ probability using $O\left(\frac{1}{\epsilon^{2}}\log\frac{1}{\delta}\right)$ queries and samples.
@@ 3 Optimizing sample access
The $L^{2}$ norm sample access arises naturally from the quantum encoding of classical vectors, which is widely utilized in the development of quantum algorithms [35, 24, 34, 37]. For a classical vector $x\in\mathbb{R}^{n}$, its corresponding quantum state $|x\rangle$ is defined as $\frac{1}{\|x\|}\sum_{i=1}^{n}x_{i}|i-1\rangle$, where the computational basis state $|i-1\rangle$ is associated with a probability proportional to $x_{i}^{2}$. However, without considering this quantum state motivation, the most natural way to assign probabilities to data elements is to make them proportional to the absolute values of the elements, rather than their squares. For a systematic comparison, we first extend Section 2 to arbitrary norms.
@@ 3.2 Inner product estimation with $\text{SQ}_{p}$
Proposition 3.
Given $\textnormal{SQ}_{p}(x)$ and $\textnormal{Q}(y)$ for $x,y\in\mathbb{R}^{n}$, the inner product $\langle x,y\rangle$ can be estimated to additive error $\epsilon||x||_{p}^{p/2}\sqrt{\langle x^{(2-p)},y^{(2)}\rangle}$ with at least $1-\delta$ probability using $O\left(\frac{1}{\epsilon^{2}}\log\frac{1}{\delta}\right)$ queries and samples.
@@
As a practical example, the quantum-inspired recommendation algorithm involves estimating the inner products between user rows [40]. Proposition 3 was evaluated on the MovieLens dataset [19], comprising 330,975 users and their ratings for 288,983 movies. The sampled pairs of rows were restricted to those with a minimum of 50 shared nonzero positions. For each pair, the constant $||x||_{p}^{p/2}\sqrt{\langle x^{(2-p)},y^{(2)}\rangle}$ in Proposition 3, which is directly related to the number of samples required for a fixed precision, was evaluated for $p=1,2$. On average, this constant was $\approx 2.16$ times smaller for $p=1$ compared to $p=2$, which implies that $\text{SQ}_{1}$ requires fewer samples than $\text{SQ}_{2}$ to achieve the same accuracy in estimating the inner product.
Assuming certain conditions, theoretical support can be provided for the experimental evidence showing that $\text{SQ}_{1}$ is preferable to $\text{SQ}_{2}$. Note that in Proposition 3, while $x$ is fixed for most of the time, the input $y$ varies across different problem instances. Therefore, it makes sense to analyze the average performance by viewing $y$ as a random variable. If we make further assumptions about the input $y$, the performance (averaged over $y$) can be expressed only in terms of $x$ and $p$, which determines the optimal value of $p$. In this respect, we now focus on deriving the optimal sample access under specific, simplified conditions.
In machine learning, processing steps like normalization are frequently applied to input data [21]. Therefore, we assume that the entries of the input vector have equal second-order moments. Even without this assumption, the law of large numbers and concentration phenomena suggest that deviations in the algorithm’s global behavior are likely negligible, since we operate on high-dimensional vectors.
Proposition 4.
$\textnormal{SQ}_{1}$ has the optimal sample efficiency, averaged over $y$, for inner product estimation in Proposition 3 if the entries of $y$ have equal second-order moments.
@@
If we are given $\textnormal{SQ}_{p}(A)$ for some matrix $A$, then estimating the value of $x^{\top}Ay$ for two inputs $x$ and $y$ with query access is also an inner product estimation. Indeed, this is a special case of trace inner product estimation introduced in the context of the quantum-inspired support vector machine [11]. Under the assumption that each $x_{i}\cdot y_{j}$ has the same second-order moment, Proposition 4 holds for this task.
@@ 3.3 Sampling from a linear combination of vectors with $\text{SQ}_{p}$
The sampling method that achieves the result of Proposition 2 utilizes rejection sampling [14, 32, 5, 42, 27]. Suppose $P(i)$ is the distribution to which we have sample access and $Q(i)$ is the target distribution. Each time we pull a sample $i$ from $P$, we compute the ratio $r_{i}:=\frac{Q(i)}{MP(i)}$ for some constant $M$. Then we output $i$ with probability $r_{i}$ and restart (or reject) otherwise.
@@
In this section, we analyze performance purely in terms of expected number of queries per sample. Other factors are discussed in Section 6. We consider the recommendation system again. Each item recommended to a user is sampled from a linear combination of a small subset of relevant user rows in the MovieLens dataset. The number of selected users ranges from 10 to 100, and the coefficients for the linear combination were drawn independently from $\mathcal{N}(0,1)$. As observed in Table 1, there is a noticeable gap in the expected number of queries per sample between $\text{SQ}_{2}$ and $\text{SQ}_{1}$. This is because the user rows are sparse and $M(p)$ behaves like $n^{p-1}$ (see Corollary 3).
Table 1: The ratio of the expected number of queries per sample when using $\text{SQ}_{2}$ to that when using $\text{SQ}_{1}$ (which is equal to $M(2)/M(1)$) for the MovieLens dataset, averaged over 1000 instances. As in Proposition 5, $n$ denotes the dimension of $x$, or equivalently, the number of users involved in the linear combination of user vectors in the context of recommendation systems.
@@
Suppose we fix the value of $p$. Then from Lemma 2, the expected number of iterations for a single valid sample, i.e., $M(p)$, behaves like $n^{p/2}$ times a constant for large $n$. Therefore, the expected query complexity for a single valid sample is $O(n^{1+p/2})$. It is evident from this expression that small $p$ is desirable. The following proposition immediately follows.
Proposition 6.
If the entries of $A$ and $x$ are i.i.d. zero mean random variables and $m,n\gg 1$, then $\textnormal{SQ}_{1}$ minimizes on average the expected number of queries per sample for sampling from a linear combination of vectors in Proposition 5, where the average is taken over $A$ and $x$.
@@
Propositions 4 and 6 collectively suggest that for the tasks of inner product estimation and sampling from a linear combination of vectors, $\textnormal{SQ}_{1}(x)$ is preferable to $\text{SQ}_{2}$ in terms of time complexities. However, it is important to note that Proposition 6 compares $\textnormal{SQ}_{1}$ with other data structures solely in terms of time complexity. If the underlying norm itself used for sampling matters, a fair comparison cannot be made. This issue is discussed further in Section 6.
@@ 4 Numerical simulations
Lemma 2 assumes $m\gg 1$ and $n\gg 1$. While the former is a reasonable assumption for quantum-inspired algorithms to have practical advantages, the latter might not be the case (consider the recommendation scenario). We plot the empirical mean of $M(p)$ for small values of $n$. The following two distributions are considered for $f$ in Lemma 2:
@@ 5 Applications to DFE
As another example, we explain how sample access affects DFE using Pauli measurements [13, 20, 45, 28]. DFE enables efficient and scalable verification of quantum states in quantum computing devices by circumventing the exponential resource requirements of full quantum state tomography [9, 13]. It is essentially a sampling-based inner product estimation algorithm. In this sense, it aligns with the principles of quantum-inspired machine learning. This prompts us to investigate whether sampling based on $L^{1}$ norm improves the performance of DFE. We provide an overview of the prerequisites and frameworks.
Suppose our goal is to prepare an $n$-qubit quantum state (with Hilbert space dimension $d=2^{n}$) in the lab to be a pure target state $\rho=|\psi\rangle\langle\psi|$, but we have instead produced an unknown state $\sigma$. It is typical to use fidelity as a metric to evaluate how similar $\sigma$ is to $\rho$ [22, 33, 43, 29, 2]. If the target state $\rho$ is pure, the fidelity between $\rho$ and $\sigma$ is defined as
@@ 5.1 Sample efficiency for the W state
In this section, we will restrict our focus to the W state [12, 4] as the target:
@@
Now suppose we sample $k$ from $\text{SQ}_{1}$, instead of $\text{SQ}_{2}$. While the assumption in Proposition 4 does not apply to the current task, it turns out that we can independently show that $\text{SQ}_{1}$ tightens the upper bound on the number of samples needed to achieve the same precision. First, select $k$ with probability $|\chi_{\rho}(k)|/Z$, where $Z\equiv\sum_{k}|\chi_{\rho}(k)|$ is the normalizing constant. Then define the estimator as
@@
[…]
@@
where $\mathbb{E}[\tilde{Y}]=Y$. Since each $M_{ij}$ is ±$\pm 1$, each sample in (4) has an absolute value of $Z/lN_{i}\sqrt{d}$. We can bound this value by upper-bounding $Z$ (see Appendix H):
@@
As in [13], one can show that
@@
Each $k_{i}$ is sampled as follows. With probability $(n-1)\sqrt{d}/2Z$, go to the second branch in (3); otherwise, go to the first branch. If we are in the first branch, sample $w$ with probability $\binom{n}{w}|n-2w|/Z^{\prime}$. The remaining steps are the same as in [13].
@@
For stabilizer states (e.g., the GHZ state [18, 30, 6]), the characteristic function yields a uniform distribution over $d$ operators. In this case, replacing $\textnormal{SQ}_{2}$ with $\textnormal{SQ}_{1}$ makes no difference.

### v3

# Toward Optimizing the Expected Performance of Sampling-Based Quantum-Inspired Algorithms

@@ 1 Introduction
Quantum algorithms have intrigued researchers for years with the potential of achieving exponential speedup compared to classical algorithms [deutsch1992rapid, shor1999polynomial, ronnow2014defining, montanaro2015quantum, kerenidis2016quantum, childs2018toward]. These algorithms leverage the principles of quantum mechanics to perform computations in parallel, potentially solving certain problems much more efficiently than classical algorithms ever could.
However, contrary to prevailing belief, it was recently shown that under certain conditions—when data is structured in a particular way—some quantum machine learning (QML) algorithms do not exhibit exponential speedup over classical ones [gilyen2018quantum, tang2019quantum, tang2021quantum, chia2022sampling]. In other words, the exponential speedup of quantum algorithms relied on the assumption that preparing the quantum state encoding of a given classical input could be achieved in sublinear time relative to the input size. This fact was first demonstrated by explicitly constructing a classical recommendation algorithm inspired by its quantum counterpart [tang2019quantum]. If classical data is preprocessed in a proper manner, a dequantized QML algorithm can be designed to run only polynomially slower than the quantum version. Specifically, the data must be prepared in a tree structure to enable efficient sampling comparable to quantum algorithms.
Since the introduction of the quantum-inspired recommendation system, several quantum algorithms have been successfully dequantized [gilyen2018quantum, tang2021quantum, chia2022sampling], highlighting both the limitation and potential of quantum computing. Under certain conditions, these algorithms are expected to achieve much better performance than known theoretical bounds, especially in extremely high-dimensional problems [arrazola2020quantum]. Therefore, it is instructive to characterize the best possible performance of these quantum-inspired algorithms for precise comparison and further development.
@@
Direct Fidelity Estimation (DFE) [flammia2011direct], a technique developed in advance of quantum-inspired algorithms, is another practical application of sampling-based inner product estimation. In Section 5, we show that for some target states, we can tighten the upper bound on the number of required measurements to perform DFE by sampling a Pauli operator based on the $L^{1}$ norm distribution for the characteristic function.
@@ 2.1 Nomenclature
[…]
@@ 2.2 $\text{SQ}_{2}$ data structure
[…]
Hereafter, the notation $\mathcal{D}_{x}$ implies that $x$ is a nonzero vector. Definition 1 has a quantum analog, since $\mathcal{D}_{x}$ is the distribution corresponding to the amplitude encoding of $x$ [schuld2018supervised]:
@@
Figure 2: BST data structures for (a) $x\in\mathbb{R}^{4}$ ($\mathrm{SQ}_{p}(x)$) and (b) $A\in\mathbb{R}^{4\times 4}$ ($\mathrm{SQ}_{p}(A)$) based on $L^{p}$ norm.
@@ 3 Optimizing sample access
The $L^{2}$ norm sample access arises naturally from the quantum encoding of classical vectors, which is widely utilized in the development of quantum algorithms [rebentrost2014quantum, kerenidis2016quantum, park2020theory, schuld2020circuit]. For a classical vector $x\in\mathbb{R}^{n}$, its corresponding quantum state $|x\rangle$ is defined as $\frac{1}{\|x\|}\sum_{i=1}^{n}x_{i}|i-1\rangle$, where the computational basis state $|i-1\rangle$ is associated with a probability proportional to $x_{i}^{2}$. However, without considering this quantum state motivation, the most natural way to assign probabilities to data elements is to make them proportional to the absolute values of the elements, rather than their squares. For a systematic comparison, we first extend Section 2 to arbitrary norms.
@@ 3.2 Inner product estimation with $\text{SQ}_{p}$
Observation 1.
[…]
@@
As a practical example, the quantum-inspired recommendation algorithm involves estimating the inner products between user rows [tang2019quantum]. Observation 1 was evaluated on the MovieLens dataset [harper2015movielens], comprising 330,975 users and their ratings for 288,983 movies. The sampled pairs of rows were restricted to those with a minimum of 50 shared nonzero positions. For each pair, the constant $x||_{p}^{p/2}\sqrt{\langle x^{(2-p)},y^{(2)}\rangle}$ in Observation 1, which is directly related to the number of samples required for a fixed precision, was evaluated for $p=1,2$. On average, this constant was $\approx 2.16$ times smaller for $p=1$ compared to $p=2$, which implies that $\text{SQ}_{1}$ requires fewer samples than $\text{SQ}_{2}$ to achieve the same accuracy in estimating the inner product.
Assuming certain conditions, theoretical support can be provided for the experimental evidence showing that $\text{SQ}_{1}$ is preferable to $\text{SQ}_{2}$. Note that in Observation 1, while $x$ is fixed for most of the time, the input $y$ varies across different problem instances. Therefore, it makes sense to analyze the average performance by viewing $y$ as a random variable. If we make further assumptions about the input $y$, the performance (averaged over $y$) can be expressed only in terms of $x$ and $p$, which determines the optimal value of $p$. In this respect, we now focus on deriving the optimal sample access under specific, simplified conditions.
In machine learning, processing steps like normalization are frequently applied to input data [ioffe2015batch]. Therefore, we assume that the entries of the input vector have equal second-order moments. Even without this assumption, the law of large numbers and concentration phenomena suggest that deviations in the algorithm’s global behavior are likely negligible, since we operate on high-dimensional vectors.
Observation 2.
$\textnormal{SQ}_{1}$ has the optimal sample efficiency, averaged over $y$, for inner product estimation in Observation 1 if the entries of $y$ have equal second-order moments.
@@
If we are given $\textnormal{SQ}_{p}(A)$ for some matrix $A$, then estimating the value of $x^{\top}Ay$ for two inputs $x$ and $y$ with query access is also an inner product estimation. Indeed, this is a special case of trace inner product estimation introduced in the context of the quantum-inspired support vector machine [ding2021quantum]. Under the assumption that each $x_{i}\cdot y_{j}$ has the same second-order moment, Observation 2 holds for this task.
@@ 3.3 Sampling from a linear combination of vectors with $\text{SQ}_{p}$
The sampling method that achieves the result of Proposition 2 utilizes rejection sampling [forsythe1972neumann, neal2003slice, casella2004generalized, thomopoulos2012essentials, legault2019accounting]. Suppose $P(i)$ is the distribution to which we have sample access and $Q(i)$ is the target distribution. Each time we pull a sample $i$ from $P$, we compute the ratio $r_{i}:=\frac{Q(i)}{MP(i)}$ for some constant $M$. Then we output $i$ with probability $r_{i}$ and restart (or reject) otherwise.
@@
In this section, we analyze performance purely in terms of expected number of queries per sample. Other factors are discussed in Section 6. We consider the recommendation system again. Each item recommended to a user is sampled from a linear combination of a small subset of relevant user rows in the MovieLens dataset. The number of selected users ranges from 10 to 100, and the coefficients for the linear combination were drawn independently from $\mathcal{N}(0,1)$. As observed in Table 1, there is a noticeable gap in the expected number of queries per sample between $\text{SQ}_{2}$ and $\text{SQ}_{1}$. This is because the user rows are sparse and $M(p)$ behaves like $n^{p-1}$ (see Observation 5).
Table 1: The ratio of the expected number of queries per sample when using $\text{SQ}_{2}$ to that when using $\text{SQ}_{1}$ (which is equal to $M(2)/M(1)$) for the MovieLens dataset, averaged over 1000 instances. As in Observation 4, $n$ denotes the dimension of $x$, or equivalently, the number of users involved in the linear combination of user vectors in the context of recommendation systems.
@@
Suppose we fix the value of $p$. Then from Observation 7, the expected number of iterations for a single valid sample, i.e., $M(p)$, behaves like $n^{p/2}$ times a constant for large $n$. Therefore, the expected query complexity for a single valid sample is $O(n^{1+p/2})$. It is evident from this expression that small $p$ is desirable. The following proposition immediately follows.
Observation 8.
If the entries of $A$ and $x$ are i.i.d. zero mean random variables and $m,n\gg 1$, then $\textnormal{SQ}_{1}$ minimizes on average the expected number of queries per sample for sampling from a linear combination of vectors in Observation 4, where the average is taken over $A$ and $x$.
@@
Observations 2 and 8 collectively suggest that for the tasks of inner product estimation and sampling from a linear combination of vectors, $\textnormal{SQ}_{1}(x)$ is preferable to $\text{SQ}_{2}$ in terms of time complexities. However, it is important to note that Observation 8 compares $\textnormal{SQ}_{1}$ with other data structures solely in terms of time complexity. If the underlying norm itself used for sampling matters, a fair comparison cannot be made. This issue is discussed further in Section 6.
@@ 4 Numerical simulations
Observation 7 assumes $m\gg 1$ and $n\gg 1$. While the former is a reasonable assumption for quantum-inspired algorithms to have practical advantages, the latter might not be the case (consider the recommendation scenario). We plot the empirical mean of $M(p)$ for small values of $n$. The following two distributions are considered for $f$ in Observation 7:
@@ 5 Applications to DFE
As another example, we explain how sample access affects DFE using Pauli measurements [flammia2011direct, huang2020predicting, zhang2021direct, leone2023nonstabilizerness]. DFE enables efficient and scalable verification of quantum states in quantum computing devices by circumventing the exponential resource requirements of full quantum state tomography [da2011practical, flammia2011direct]. It is essentially a sampling-based inner product estimation algorithm. In this sense, it aligns with the principles of quantum-inspired machine learning. This prompts us to investigate whether sampling based on $L^{1}$ norm improves the performance of DFE. We provide an overview of the prerequisites and frameworks.
Suppose our goal is to prepare an $n$-qubit quantum state (with Hilbert space dimension $d=2^{n}$) in the lab to be a pure target state $\rho=|\psi\rangle\langle\psi|$, but we have instead produced an unknown state $\sigma$. It is typical to use fidelity as a metric to evaluate how similar $\sigma$ is to $\rho$ [jozsa1994fidelity, nielsen2010quantum, wilde2013quantum, liang2019quantum, baldwin2023efficiently]. If the target state $\rho$ is pure, the fidelity between $\rho$ and $\sigma$ is defined as
@@ 5.1 Sample efficiency for the W state
In this section, we will restrict our focus to the W state [dur2000three, cabello2002bell] as the target:
@@
Now suppose we sample $k$ from $\text{SQ}_{1}$, instead of $\text{SQ}_{2}$. While the assumption in Observation 2 does not apply to the current task, it turns out that we can independently show that $\text{SQ}_{1}$ tightens the upper bound on the number of samples needed to achieve the same precision. First, select $k$ with probability $|\chi_{\rho}(k)|/Z$, where $Z\equiv\sum_{k}|\chi_{\rho}(k)|$ is the normalizing constant. Then define the estimator as
@@
Each Pauli operator $W$ can be represented by two binary vectors $\textbf{j},\textbf{k}\in\{0,1\}^{n}$ which denote the positions for $X$ and $Z$, respectively (up to a global phase). For example, $W=\mathbb{I}\otimes X\otimes Y\otimes Z$ corresponds to $\textbf{j}=(0,1,1,0)$ and $\textbf{k}=(0,0,1,1)$ (note that $Y=iXZ$). Sampling $k$ from $\text{SQ}_{1}$ is equivalent to sampling $\textbf{j},\textbf{k}$ from the following distribution, which directly follows from flammia2011direct:
@@
where $\mathbb{E}[\tilde{Y}]=Y$. Since each $M_{ij}$ is $\pm 1$, each sample in (4) has an absolute value of $Z/lN_{i}\sqrt{d}$. We can bound this value by upper-bounding $Z$ (see Appendix H):
@@
As in [flammia2011direct], one can show that
@@
Each $k_{i}$ is sampled as follows. With probability $(n-1)\sqrt{d}/2Z$, go to the second branch in (3); otherwise, go to the first branch. If we are in the first branch, sample $w$ with probability $\binom{n}{w}|n-2w|/Z^{\prime}$. The remaining steps are the same as in [flammia2011direct].
@@
For stabilizer states (e.g., the GHZ state [greenberger1989going, mermin1990quantum, caves2002unknown]), the characteristic function yields a uniform distribution over $d$ operators. In this case, replacing $\textnormal{SQ}_{2}$ with $\textnormal{SQ}_{1}$ makes no difference.